### v6

# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

Abstract
Online Multi-Object Tracking (MOT) plays a pivotal role in autonomous systems. The state-of-the-art approaches usually employ a tracking-by-detection method, and data association plays a critical role. This paper proposes a learning and graph-optimized (LEGO) modular tracker to improve data association performance in the existing literature. The proposed LEGO tracker integrates graph optimization, which efficiently formulates the association score map, facilitating the accurate and efficient matching of objects across time frames. To further enhance the state update process, the Kalman filter is added to ensure consistent tracking by incorporating temporal coherence in the object states to further enhance the state update process. Our proposed method, utilising LiDAR alone, has shown exceptional performance compared to other online tracking approaches, including LiDAR-based and LiDAR-camera fusion-based methods. LEGO ranked $3^{rd}$ among all trackers (both online and offline) and $2^{nd}$ among all online trackers in the KITTI MOT benchmark for cars [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking.php, at the time of submitting results to KITTI object tracking evaluation ranking board. Moreover, our method also achieves competitive performance on the Waymo open dataset benchmark.
Index Terms:
Multi-object tracking, online tracking, transformer, graph optimization, graph neural network, data association, track management, LiDAR, point cloud, autonomous driving
I Introduction
Tracking is a crucial technology utilized in various systems within the field of intelligent transportation systems (ITS), like pedestrian tracking [1, 2], drone monitoring [3, 4], target tracking [5], traffic monitoring [6, 7], Advanced Driver Assistance Systems (ADAS) and Autonomous Driving (AD) [8, 9]. To develop robust and accurate tracking systems effectively, various approaches have been investigated by using different sensor modalities, encompassing 2D multi-ojbect-tracking (MOT) methods using camera images, 3D MOT techniques relying solely on LiDAR, and 3D MOT methods that leverage both camera and LiDAR. In the realm of 2D MOT, notable methods, such as FairMOT [10], ByteTrack [11], and StrongSort [12] predominantly employ camera data for tracking objects. Although these methods provide valuable insights, their performance is constrained by the limitations of working with 2D representations alone. On the other hand, 3D MOT techniques exclusively employ LiDAR to capture 3D information about the tracked objects. AB3DMOT [21] serves as an exemplary method within this category, demonstrating the efficacy of LiDAR-based approaches for precise object localization and spatial understanding. Furthermore, there are 3D MOT methods that exploit the synergies between camera and LiDAR systems. Probabilistic CBMOT [48] represents one such approach, where both sensor modalities are utilized to improve tracking performance. By combining the strengths of cameras and LiDAR sensors, these methods aim to achieve enhanced object tracking accuracy. A majority of these MOT methods can be classified as tracking-by-detection methods, in which various 3D object detectors, e.g., PointRCNN [13], PointGNN [14], CenterPoint [15], and CasA [16], etc, have been employed to provide estimated bounding boxes as measurements input to the MOT pipeline. Some work, such as [3, 4], may use drone monitoring to collect 2D video image information and perform the tracking based on a temporal transformer. As drone video is usually two-dimensional, it lacks the rich depth information available from other sources. However, continuous point cloud data is a form of 3D video for real-world representation. This is quite different from conventional 2D video, as it captures spatial structure and depth information directly in three dimensions, enabling more accurate modeling, understanding, and tracking of dynamic scenes in complex environments.
Despite many efforts have been made in this area, MOT systems still suffer from erroneous measurements generated from object detector and inaccurate data association. Specifically, erroneous measurements from object detectors lead to misaligned or imprecise bounding boxes, which degrade the quality of input data for tracking. Inaccurate data association, on the other hand, results in track identity switches or fragmented tracks, making it difficult to maintain consistent object trajectories. Consequently, these issues significantly hinder the overall performance of MOT systems.
To address these challenges, we propose a novel framework that explicitly tackles both issues through the introduction of two key components: the offset correction module and the Adjacency Matrix guided Graph Network (AMGN) score calculation module. The offset correction module is designed to correct the erroneous bounding boxes generated by the detector, resulting in more accurate and reliable object localization. Meanwhile, the AMGN score calculation module redefines the data association problem as a bipartite graph matching problem, leveraging a combination of graph structure information, an LSTM, and a new self-attention structure whose coefficients is derived from the learnable adjacency matrix to enhance relational reasoning and reduce information redundancy. This integration allows the Graph Neural Network (GNN) to effectively capture complex dependencies between measurements and predicted tracks, improving the accuracy of data association. Lastly, by integrating with Kalman filter, our approach facilitates more robust state updates and improves overall tracking performance.
The contributions of this paper are outlined as follows:
  • A easily comprehensible and tunable LEarning and Graph-Optimized (LEGO) modular tracker is proposed for online MOT. LEGO introduces two learning modules as its core, an offset correction module and a Adjacency Matrix guided Graph Network (AMGN) score calculation module.
  • The offset correction module is designed to predict the offset between these detection results and the ground truth. The primary purpose of this mechanism is to rectify prediction outcomes, thereby improving the accuracy of the model’s predictive capabilities. The AMGN score calculation module which contains LSTM and $K$ AMGN blocks, leverages the LSTM and graph structure information to perform message passing and update node features using the self-attention coefficients derived from a learnable adjacency matrix. By doing so, the AMGN score calculation module can directly computes the association score.
  • Empirical evaluation and performance analysis of the proposed offset correction and AMGN score calculation modules, demonstrate their effectiveness in improving tracking performance, as well as the proposed LEGO tracker’s state-of-the-art performance in both the KITTI MOT benchmark and Waymo 3D MOT benchmark.
The structure of this paper is organized as follows: Section II presents a thorough review of related works in the field of LiDAR-based MOT within the context of autonomous driving applications. Section III introduces the proposed LEGO Modular tracker, elucidating its key components and functionalities. Subsequently, in Section IV, the experimental results are presented and analyzed in detail. Finally, the paper concludes with Section V, summarizing the key findings and contributions of the study.
II Related Work
II-A MOT with LiDAR Only
II-A1 Model-based Methods
Similar to MOT tasks in image filed [17][18], contemporary tracking systems in ADAS and AD applications often employ global nearest neighbor methods and heuristics for data association. In LiDAR-based MOT, several methods have emerged that rely exclusively on LiDAR sensors. Chiu et al. [23] made a pioneering contribution by integrating Mahalanobis distance with AB3DMOT, establishing a benchmark for addressing LiDAR-based 3D MOT challenges. Similarly, SimpleTrack [28] introduced a generalized version of 3D IoU, known as GIoU, as the scoring mechanism for tracking-by-detection tasks. Bytetrackv2 [38] employed a hierarchical data association strategy to identify genuine objects within low-score detection boxes, effectively mitigating issues related to object loss and fragmented trajectories. Additionally, this system employed Non-Maximum Suppression (NMS) to preprocess object detections. Empirical evidence demonstrated that combining GIoU with NMS preprocessing enhanced overall tracking performance.
Maintaining tracks even when objects are no longer visible is a common feature shared by ImmortalTracker [29] and PC3T [45]. This helps reduce identification switches and fragmented tracks. ACK3DMOT [39] introduced a cost matrix for tracking-by-detection tasks based on a joint probability function that considers appearance, geometry, and distance correlation between detected bounding boxes and predicted objects. When combined with an adaptive cubature Kalman filter, this approach achieved enhanced tracking performance. PF-MOT [32] utilized a cluster-based earth-mover distance, Euclidean distance, and feature similarity to construct the cost matrix. Considering uncertainties, UG3DMOT [43] evaluated data association based on random vectors, where the similarity between two multidimensional distributions was evaluated using the Jensen-Shannon divergence.
As the raise of applying random finite set (RFS) in MOT applications [8], there are alternative algorithms where each potential object is modelled using a Bernoulli process with probabilistic object existence. RFS-M3 [24] utilized the Poisson multi-Bernoulli mixture filter based on a random finite set (RFS) to tackle the LiDAR-based MOT problem. Through systematic comparative analysis, GNN-PMB [34] demonstrated that the contemporary RFS-based Bayesian tracking framework outperformed the traditional random vector-based Bayesian tracking framework. BP-Tracker [30] presented a factor graph formulation of the MOT problem and employed a belief propagation algorithm to compute the marginal association probability, representing a significant advancement in the field.
II-A2 Deep-learning based Methods
SimTrack [25] and CenterTube [41] introduce end-to-end trainable models for joint detection and tracking, leveraging raw point cloud as input. OGR3MOT [26], Batch3DMOT [27], PolarMOT [42], Rethinking3DMOT propose sophisticated graph structures based on neural message passing, enabling online execution of detection and tracking processes. The ENBP-Tracker [33] integrates GNN into its design, distinguishing itself by combining the network with a belief propagation tracker. This hybrid tracking architecture enhances the robustness and efficacy of the tracking process, showcasing the potential of merging traditional tracking methods with advanced neural network architectures. Intertrack [35] TransMOT [31] adopt transformer structures to generate discriminative object representations for data association. Minkowski-Tracker [36], PC-TCNN [46] and ShaSTA [37], employ proposal networks to extract features from various feature maps, facilitating the learning of affinity matrix for point cloud-based MOT.
II-B MOT with LiDAR and Camera Fusion
The effectiveness of MOT can be further enhanced by adopting a fusion approach that combines data from LiDAR and camera sources. Several methods, such as Probabilistic3DMM [47], CBMOT [48], GNN3DMOT [52], and MF-Net [49] have utilized this fusion approach to leverage the complementary strengths of different sensor modalities, leading to more comprehensive and precise object tracking. IMSF MOT [57] proposes a novel feature fusion method using Pointnet++ to extract more discriminative features and improve multiple object tracking performance.
DeepFusionMOT [61] and EagerMOT [58] improve tracking performance compared to Probabilistic3DMM by implementing a two-stage data association scheme. This scheme leverages 3D detection data obtained from LiDAR and camera inputs, and 2D detection data obtained solely from the camera. By combining information from multiple modalities, a more comprehensive understanding of the environment is achieved, potentially enhancing the robustness and accuracy of the tracking process. Building upon EagerMOT, AlphaTrack [50] introduces a feature extractor that concatenates image and point cloud information as input to enhance performance. Additionally, other relevant works, such as DualTracker [60], HIDMOT [64], CAMO-MOT [9], MSA-MOT [65], and JMODT [59] utilize PointGNN or PointRCNN as their 3D object detectors to estimate object detections. They are combined with detection proposals from image data, and a hybrid multi-modal input is employed for the association within the tracking-by-detection task, potentially improving the overall performance and accuracy of the tracking system.
II-C MOT with Graph Neural Network based Methods
In recent years, GNN-based tracking methods have attracted considerable research interest, primarily due to their capability to model Multi-Object Tracking (MOT) as a graph structure, allowing object nodes to effectively exchange and aggregate features. GNN3DMOT [52] was among the first to propose a GNN-based 3D MOT framework, integrating both 2D and 3D feature learning. Their method employs a feature extractor to obtain motion and appearance features from both 2D images and 3D point clouds, subsequently feeding these features into a GNN to facilitate feature interaction among objects. Furthermore, [51] introduced an MOT approach comprising two separate graph networks: an appearance graph network and a motion graph network. These two GNNs independently compute similarity measures between detected objects and existing trackers through a four-step updating module. The final affinity matrix is then derived by combining these two similarity scores through a weighted fusion mechanism. Additionally, [53] proposed an iterative Graph Convolutional Network (GCN) clustering method designed to reduce computational complexity. This method ranks generated proposals based on estimated quality scores, effectively maintaining proposal quality while significantly enhancing computational efficiency. Similarly, [54] presented a GCN-based MOT approach, modeling the relationships between existing tracklets and intra-frame detections as a general undirected graph to improve tracking robustness and accuracy. Most existing methods perform message passing and updating with Graph Attention Networks (GAT) to ensure that different nodes contribute differently during the message passing process. However, GAT handles message passing and node updating on a node-by-node basis, which overlooks the global structure of the graph. This limitation may result in the loss of critical information embedded within the overall graph structure. To address this issue, we propose the AMGN block, which incorporates a learnable adjacency matrix into the self-attention mechanism during the message passing and updating process. This design enables the model to capture both local and global structural information, leading to more effective graph representation.
III The Proposed LEGO Modular Tracker
Figure 1: The proposed LEGO tracker seamlessly integrates several modules. First, the offset correction module refines detection results and corrects detected errors. Simultaneously, the motion prediction module utilizes the tracked state $\bm{X}_{t}$ to forecast the next state $\bm{X}_{t|t-1}$. The feature extraction module then derives 3D feature maps for both predicted and detected objects. Using this information, the AMGN score calculation module constructs a bipartite graph and computes the AMGN score matrix. The association module subsequently integrates the geometry and motion cost matrix with the AMGN score matrix, yielding an association map. Finally, the state update module refines the predicted state, and the track management module processes the matched detected states, completing the tracking cycle.
Within this section, we will introduce the LEGO modular tracker. Firstly, we will present its overall framework, followed by a detailed elaboration of each module. Fig. 1 shows the proposed LEGO tracker, which comprises several modules: motion prediction, proposed offset correction, feature extraction, proposed AMGN score calculation, association, state update, and track management.
III-A The Overview Framework of LEGO Modular Tracker
Initiating the process, the 3D point cloud is fed into a 3D object detector. To refine these detection results, an offset correction module is introduced. This module predicts the offset between the detection outputs and the ground truth, smoothing the detector’s predictions. Simultaneously, a motion prediction module uses the tracked state from the previous time step to predict the current state of each object, by using Constant Acceleration (CA) motion model [22].
Next, the outputs from the motion prediction and offset correction modules, along with the 3D point cloud, are passed through a feature extraction module. This module, based on the PointNet architecture [73], generates robust 3D feature maps. Using these feature maps, a bipartite graph $\bm{G}=\langle\bm{N},\bm{E}\rangle$ is constructed, where $\bm{N}$ represents nodes for detected and predicted objects, and $\bm{E}$ represents potential associations between them. The bipartite graph $\bm{G}$ and a learnable adjacency matrix are then input into the AMGN block under the proposed AMGN score calculation module, which outputs the AMGN score matrix $\bm{B}$. Using LSTM and $K$ AMGN blocks with the learnable adjacency matrix, the AMGN score calculation module captures and refines the structural relationships within the graph, quantifying the relational strengths between nodes. The final association cost matrix is derived by combining the AMGN score matrix $\bm{B}$, with the geometry and motion cost matrix $\bm{A}$ which is calculated by following PC3T [45].
The data association module then matches detected and predicted objects based on the association cost matrix. Finally, the resulting association map is passed to the state update module, which updates object states using the standard Kalman filter update. The track management module handles unmatched detected states and predicted states by using classical $M/N$ logic. Specifically, if a predicted state remains unmatched for more than $N_{t}$ consecutive frames, it is removed as leaving the scene. Otherwise, unmatched predicted states are retained, assuming the object may reappear in subsequent frames. The track management module also consider all unmatched objects as potential new objects entering the scene. In order to avoid creating false positive trajectories, new trajectory will not be created for the unmatched detection until it has been continually matched in the next $M_{t}$ frames.
III-B Motion Prediction Module
Certain MOT techniques prefer to use a constant-velocity motion model based on the Kalman Filter. This model assumes that the object moves at a nearly constant speed, which may not hold true in several real-world situations. As a result, consecutive missed detections can lead to considerable errors in motion prediction. To overcome this problem, the constant acceleration (CA) motion model [22] is used based on the Kalman Filter, which offers a more precise representation of the object state. In the CA motion model, the object state $\bm{X}_{t-1}$ at a specific temporal instance t1t-1 is represented as $\left[\bm{pos}_{t-1},\bm{v}_{t-1},\bm{\alpha}_{t-1}\right]^{T}$, where $\bm{pos}_{t-1}$ denotes the position, $\bm{v}_{t-1}$ represents the velocity, and $\bm{\alpha}_{t-1}$ indicates the orientation of the object. The prediction of the mean state $\bm{X}_{t|t-1}$ and covariance $\bm{P}_{t|t-1}$ using Kalman prediction equations with the CA motion model is as follows:
$\bm{X}_{t|t-1}=\bm{S}_{A}\bm{X}_{t-1},$ (1)
$\bm{P}_{t|t-1}=\bm{S}_{A}\bm{P}_{t-1}\bm{S}_{A}^{T}+\bm{Q},$ (2)
where $\bm{S}_{A}$ is the state transition matrix, $\bm{Q}$ represents the motion noise covariance matrix, $\bm{I}$ denotes the identity matrix, $\bm{O}$ represents the zero matrix, $n$ denotes the state dimension, and $\delta$ and $\mathbf{a}$ are sensor-related hyper-parameters.
Figure 2: The structure of the offset correction module. The detection result, which comprises the parameters $x_{p}^{d},y_{p}^{d},z_{p}^{d},w^{d},h^{d},l^{d}$, is input into the offset prediction head. The offset prediction head can be constructed” as a simple neural network, which computes the offset between the ground truth and the obtained detection result.
III-C Offset Correction Module
Our investigation has revealed a significant issue related to the accuracy of the state inferred by the detector, especially when errors arise from the 3D object detector. In order to mitigate this potential source of inaccuracies, an offset correction module is proposed. This module serves the crucial purpose of rectifying the detection outcomes generated by the 3D object detector, thereby enhancing the overall accuracy of the system. The architecture and components of this offset correction module are depicted in Fig. 2. This module comprises a 3D convolution layer, batch normalization layer, and Multi-Layer Perception (MLP) with the residual connection. The offset obtained from the head is $O_{pred}$, and the ground truth offset is:
$O_{gt}=\log\left(\frac{gt-dr}{\beta}\right),$ (3)
where $gt=\left(x_{p}^{g},y_{p}^{g},z_{p}^{g},w^{g},h^{g},l^{g}\right)$} represents the ground truth bounding box coordinates $(x_{p}^{g},y_{p}^{g},z_{p}^{g})$, width $w^{g}$, height $h^{g}$, length $l^{g}$; $dr=\left(x_{p}^{d},y_{p}^{d},z_{p}^{d},w^{d},h^{d},l^{d}\right)$} indicates the object detection bounding box extracted from the 3D object detector, and $\beta$ is the scaling factor.
The offset correction module is directly supervised by the ground-truth offset. The Hungarian algorithm is used to match the ground truth bounding box and the predicted bounding box. Specifically, the loss function is defined for training the offset correction module as the Smooth L1 loss between the predicted offset $O_{pred}$ and the ground-truth offset $O_{gt}$, formulated as follows:
$L_{offset}=\text{SMOOTHL1}\left(O_{pred},O_{gt}\right).$ (4)
The ground truth data is utilized exclusively in the training phase of the offset correction module. Specifically, the module is trained to predict the offset between the ground truth annotations and the detector outputs, allowing it to learn systematic error patterns inherent in the detector. However, at inference time, the ground truth data is not available. Instead, the trained offset correction module directly predicts and rectifies the offsets based solely on the detection results provided by the 3D object detector.
III-D Feature Extraction Module
The feature extraction module uses a combination of information from two important parts: the detected state after the offset correction module and the predicted state from the motion prediction module. This combination helps create feature maps that highlight important characteristics of the object being targeted, using data from the 3D point cloud within the bounding box.
Specifically, the 3D point clouds within the detected and predicted 3D bounding boxes are cropped, feed into a PointNet encoder [73] to generate corresponding feature maps, denoted as $\bm{F}^{det}_{3d}=\left(\bm{f}^{det}_{3d_{1}},\bm{f}^{det}_{3d_{2}},\dots,\bm{f}^{det}_{3d_{J}}\right)$ and $\bm{F}^{pred}_{3d}=\left(\bm{f}^{pred}_{3d_{1}},\bm{f}^{pred}_{3d_{2}},\dots,\bm{f}^{pred}_{3d_{U}}\right)$. Sequentially, the 3D point cloud feature maps are fed into the MLP-based embedding network, to extract embedded features. This process can be described as
$\mathbf{F}=\mathtt{MLP}\left(\bm{F}^{det}_{3d}\right),$ (5)
$\mathbf{F^{*}}=\mathtt{MLP}\left(\bm{F}^{pred}_{3d}\right),$ (6)
where $\mathbf{F}$ and $\mathbf{F^{*}}$ denote the output feature maps of the feature extraction module.
III-E AMGN Score Calculation Module
Utilizing the feature map extracted from the feature extraction module, an AMGN score matrix, denoted as $\bm{B}$, is calculated through the AMGN score calculation module. Like [47], the purpose of employing such a matrix is to measure the similarity between different detections and predictions. This section details the AMGN Score Calculation Module in three parts: graph construction, AMGN blocks, and the loss function.
III-E1 Graph Construction
Most recent works, which applied GNN for 3D MOT, e.g., PTP [67], Rethinking3DMOT [44], 3Dmotformer [68], use a fully connected graph with the dimension of $\left(J+U\right)\times\left(J+U\right)$ where $J$ and $U$ denote the number of detected nodes and the number of predicted nodes in the graph, respectively. However, such a graph entails redundant computations due to the simultaneous inclusion of detection and prediction nodes.
To this end, following the approach of GNN3DMOT [52], the prior knowledge that matching should only occur across frames is utilized, and a bipartite graph is constructed. The structure of the graph is represented as a matrix of dimensions $J\times U$. The bipartite graph mitigates this computational redundancy, effectively reducing the dimension of the association score matrix. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with the nodes $\bm{N}$ are split into two distinct sets: one for detected objects $\bm{N}_{det}$ and one for predicted objects $\bm{N}_{pre}$. The set $\bm{E}$ of edges captures the associations between the nodes.
Figure 3: AMGN score calculation module. In the initial stage, a bipartite graph is fed into the multi-head self-attention (MHSA) block along with an LSTM block to extract features. This data, together with the initialization of the learnable adjacency matrix, are fed into the AMGN block. The AMGN block performs node message passing and updating using the self-attention coefficients derived from the learnable adjacency matrix. Subsequently, the learnable adjacency matrix is updated with the new feature representation obtained after the message passing and updating process. The new feature representation and updated learnable adjacency matrix are then fed into the next AMGN block. The final output is the AMGN score matrix $\bm{B}$, which is the latest learnable adjacency matrix from the last AMGN block.
III-E2 AMGN Blocks
GNN3DMOT [52] leverages the Graph Attention Network (GAT) [75] to enable different contributions from individual nodes during message passing. However, GAT processes nodes independently, and thus it may fail to capture important contextual information embedded in the global graph structure.
To overcome this, the AMGN block is proposed, which introduces a learnable adjacency matrix into the self-attention calculation. This matrix contains the structural information of the graph, allowing the model to continuously refine its understanding of the graph structure, adapt to the evolving feature representations of the nodes, and learn the strength of associations between any pair of nodes. This approach ensures that the attention mechanism incorporates both the node features and the global structural information of the graph, as captured by the learnable adjacency matrix.
The module’s process begins by generating robust initial embeddings, which are then iteratively refined by K AMGN blocks. First, a Multi-Head Self-Attention (MHSA) encoder extracts preliminary features from the concatenated detected and predicted feature maps. An LSTM network then processes these features to enhance temporal coherence, generating refined initial embeddings:
$\mathbf{Z}^{0}=\mathtt{LSTM}(\mathtt{MHSA}([\mathbf{F},\mathbf{F}^{*}])),$ (7)
where $\bm{Z^{0}}$ is the bipartite graph with the embedding representation sets of detected nodes and predicted nodes that contains $\bm{z}^{0}_{0},\dots,\bm{z}^{0}_{j},\dots,\bm{z}^{0}_{J}]$ and $\bm{z}^{*0}_{0},\dots,\bm{z}^{*0}_{u},\dots,\bm{z}^{*0}_{U}]$.
After obtaining the initial embeddings, the K AMGN blocks iteratively refine these embeddings and update the learnable adjacency matrix as shown in Fig. 3. The learnable adjacency matrix is initialized as $\bm{B}^{0}$, where its element $B^{0}_{j,u}$ between the $j$-th node in the set of nodes for detected objects and the $u$-th node in the set of nodes for predicted objects.
For the k-th AMGN block, a raw attention coefficient $e^{k}_{j,u}$ between $j$-th detected node and $u$-th predicted node is first calculated as:
$e^{k}_{j,u}=\mathtt{LeakyReLU}(a^{T}[w^{*}_{u}z^{*k-1}_{u}||w_{j}z^{k-1}_{j}]),$ (8)
where $a$ is the learnable weight vector, $||$ denotes concatenation, and $w^{*}_{u},w_{j}$ are elements in a learnable weight matrix.
Different from GAT, these attention coefficients are modified by combining them with the learnable adjacency matrix $\bm{B}^{k-1}$ from the previous block as shown:
$e^{’k}_{j,u}=e^{k}_{j,u}B^{k-1}_{j,u},$ (9)
where $B^{k-1}_{j,u}$ is the association score between $j$-th detected node and $u$-th predicted node from the prior step.
The modified coefficients are then normalized using a softmax function as:
$\alpha^{k}_{j,u}=\frac{\mathrm{exp}(e^{’k}_{j,u})}{\sum\mathrm{exp}(e^{’k}_{j,u})}.$ (10)
With these normalized attention coefficients, the node embedding is updated by aggregating neighbour information. The updated embedding for $j$-th detected node is calculated as:
$z^{’k}_{j}=\mathtt{LeakyReLU}(\sum a_{j,u}w_{u}{z^{*k}_{u}}).$ (11)
We do the same for the $u$-th predicted node to get the updated feature representation as $z^{’*k}_{u}$.
Finally, the value in the learnable adjacency matrix $\bm{B}^{k}$ is updated as shown:
$B^{k}_{j,u}=\mathtt{LeakyReLU}(\mathbf{b}^{T}[z^{’k}_{j}||z^{’*k}_{u}]),$ (12)
where $\mathbf{b}^{T}$ represents the learnable weight vector at the $k$-th module. The output of the $k$-th AMGN block serves as the input to the subsequent $(k+1)$-th AMGN block. Ultimately, the AMGN score matrix $\bm{B}$ is defined as $\bm{B}^{K}$, which is the output of the final AMGN block.
III-E3 Loss Function
The corresponding loss function for training the network to generate AMGN score matrix $\bm{B}$, association matrix loss, is expressed mathematically as
$L_{a}=\sum_{j}^{J}\sum_{u}^{U}\left[-y_{j,u}\log B_{j,u}-\left(1-y_{j,u}\right)\log\left(1-B_{j,u}\right)\right].$ (13)
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware that $y_{j,u}$ is a binary indicator, which is either 0 or $1$, as it represents the ground truth of the match between the $u$-th predicted node and the $j$-th detected node. When there is a clear match, $y_{j,u}$ is assigned as 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, a match is determined based on the continuity of object identity across time frames. In the match, a pair is considered as matched if the closest ground-truth box to the tracking box in the previous time frame and the closest ground-truth box to the detection box in the current time frame have the same index and their IoU with the closest ground-truth box is more than 0.7. The value of $B_{j,u}$ is obtained from the AMGN score matrix $\bm{B}$ and ranges between 0 and 1.
III-F Association Module
In 3D MOT, the data association module plays a crucial role in determining the correspondence between detected and predicted objects, to update the state of the objects. Most of the methods used in 3D MOT rely on the minimization of the geometry association cost matrix. This is done by evaluating the overlap or distance between the 3D bounding boxes of the predicted and detected objects, using the global nearest neighbor principle and typically the Hungarian algorithm [76]. Recently, [45] proposed to incorporate geometry, motion and appearance features from images to construct such association cost matrix. In our proposed approach, we follow a similar idea to define the geometry and motion cost matrix, but using AMGN score matrix $\bm{B}$ to replace the appearance features. The details are discussed in the follow parts.
Same as [45], the combination of geometry and motion cost matrix is expressed by the mathematical formulation
$\bm{A}=\bm{Ge}+\bm{Mo},$ (14)
where the matrix $\bm{Ge}$ denotes a geometry similarity matrix. This $\bm{Ge}$ matrix establishes the correlations between the detected state after the offset correction module, denoted as $\bm{Y}_{t}^{cor}$, and the predicted state $\bm{X}_{t|t-1}$, via:
$\bm{Ge}=\mathtt{IoU}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)+\mathtt{CEN}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right).$ (15)
The $\mathtt{IoU}$ in Eq. (15) is the cost from 3D IoU, which can be computed as
$\mathtt{IoU}=\frac{\mathtt{VI}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)}{\mathtt{VOL}\left(\bm{Y}_{t}^{cor}\right)+\mathtt{VOL}\left(\bm{X}_{t|t-1}\right)-\mathtt{VI}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)},$ (16)
where $\mathtt{VOL}$ is the volume calculated by $w,h,l$ in the state and $\mathtt{VI}$ denotes the volume intersection. The $\mathtt{CEN}$ in Eq. (15) is the cost from the centroid (in 3D Cartesian coordinate) [45] which can be computed as
[…]
where $\bm{p}$ refers to the global coordinates ($x_{p},y_{p},z_{p}$) for detected state (after the offset correction module) $Y_{t}^{cor}$ and $\bm{\hat{p}}$ is the corresponding value of predicted state $X_{t|t-1}$.
[…]
where $w_{ang}$ is the weight for angle difference, $w_{velo}$ is the weight for velocity difference, whereas $\tilde{v}$ and $\hat{v}$ denote the velocities of the predicted object states and detected object states respectively.
However, the current approach to calculating the matrix $\bm{A}$ is limited to considering only geometry and motion costs, neglecting the valuable information embedded in the 3D features present within the 3D point cloud. To address this limitation, we propose the final association cost matrix $\bm{C}$, which incorporates the geometry and motion cost matrix $\bm{A}$ and the AMGN score matrix $\bm{B}$, respectively. This amalgamation is achieved through a weighted linear combination of $\bm{A}$ and $\bm{B}$:
$\bm{C}=\bm{A}-w_{B}\bm{B},$ (19)
where $w_{B}$ represents the weight assigned to the AMGN score matrix $\bm{B}$. Notably, the subtraction operation between matrix $\bm{A}$ and $\bm{B}$ in Eq. (19) is performed due to the score being determined as a negative cost. By incorporating the AMGN score matrix alongside the geometry and motion cost matrix, our proposed method enriches the data association process, allowing for a more comprehensive evaluation of 3D features. With the final association cost matrix $\bm{C}$, as input, the Hungarian algorithm is used to solve the 2D assignment problem, identifying the matched pairs and unmatched objects.
IV Experiments and Analysis
In this section, we discuss the results of our experiment. We will begin by providing an overview of the dataset used, followed by a description of the implementation details. Next, we will present a thorough ablation study. Lastly, we will compare our results with those of the baseline models.
TABLE I: Comparison of the proposed method and other state-of-the-art LiDAR-only trackers on front view 2D MOT tracking results using KITTI car test dataset.
[…]
AB3DMOT (IEEE IROS 2020)* [21] L 69.99 69.33 86.85 83.61 85.23 66.92 113 206
PC3T (IEEE T-ITS 2022)*# [45] L 77.80 81.59 86.07 88.81 84.26 80.00 225 201
Batch3DMOT (IEEE RA-L 2022)* [27] L $N/A$ $N/A$ $N/A$ 88.60 86.80 76.70 19 74
LEGO (Ours)* L 78.05 79.22 88.08 88.97 86.92 80.92 286 71
PolarMOT (ECCV 2022)** [42] L 75.16 76.95 87.12 85.08 85.63 80.92 462 599
CenterTube (IEEE T-MM 2023)*** [41] L 71.25 69.24 86.85 86.97 85.19 78.46 191 344
LEGO (Ours)** L 79.52 83.34 87.49 88.14 86.06 87.54 290 117
UG3DMOT (Signal Processing 2024)**** [43] L 78.60 82.28 87.84 87.98 86.56 79.08 30 360
LEGO (Ours)**** L 80.75 83.27 87.92 90.61 86.66 87.85 214 109
    The metrics are reported by using PointRCNN [13] as 3D object detector.
    The metrics are reported by using PointGNN [14] as 3D object detector.
    The metrics are reported by using CenterPoint [15] as 3D object detector, which has much better detection performance than PointGNN in general.
    The metrics are reported by using CasA [16] as 3D object detector.
    Note: Since the reported metrics of PC3T are based on the trajectory refinement, which makes PC3T an offline smoother rather than online tracker like all others, thus actual performance of online tracking version of PC3T is lower.
TABLE II: Comparison of the proposed method and other state-of-the-art LiDAR and camera fusion-based trackers on front view 2D MOT tracking results using KITTI car test dataset.
[…]
JMODT (IEEE IROS 2021) [59] C+L 70.73 68.76 86.95 85.35 85.37 77.39 350 693
DeepFusionMOT (IEEE RA-L 2022)* [61] C+L 75.46 80.05 86.70 84.63 85.02 68.61 84 472
StrongFusionMOT (IEEE S-J 2022)* [62] C+L 75.65 79.84 86.74 85.53 85.07 66.15 58 416
Feng et al., (IEEE T-IV 2024)* [63] C+L 74.81 $N/A$ $N/A$ 84.82 85.17 $N/A$ $N/A$ $N/A$
LEGO (Ours)* L 78.05 79.22 88.08 88.97 86.92 80.92 286 71
EagerMOT (IEEE ICRA 2021)** [58] C+L 74.39 74.16 87.17 87.82 85.69 76.15 239 390
MSA-MOT (Sensors 2022)** [65] C+L 78.52 82.56 87.00 88.01 85.45 86.77 91 428
IMSF MOT (IEEE T-ITS 2023) [57] C+L 72.44 68.02 $N/A$ 90.32 85.47 86.46 526 270
DualTracker (IEEE T-IV 2023)** [60] C+L 74.24 $N/A$ $N/A$ 88.05 85.6 80.77 148 $N/A$
HIDMOT (IEEE T-VT 2023)** [64] C+L 75.90 77.22 $N/A$ $N/A$ $N/A$ $N/A$ $N/A$ $N/A$
LEGO (Ours)** L 79.52 83.34 87.49 88.14 86.06 87.54 290 117
MMF-JDT (IEEE RA-L 2025)*** [69] C+L 79.52 84.01 $N/A$ 88.06 86.24 $N/A$ 37 $N/A$
LEGO (Ours)*** L 80.75 83.27 87.92 90.61 86.66 87.85 214 109
    The metrics are reported by using PointRCNN [13] as 3D object detector.
    The metrics are reported by using PointGNN [14] as 3D object detector.
    The metrics are reported by using CasA [16] as 3D object detector.
TABLE III: 3D MOT tracking results of the proposed method and other state-of-the-art trackers on KITTI car validation dataset, by following the evaluation protocol in [21].
[…]
AdaptiveNoiseCov (IEEE T-IV 2024)+ [40] L 93.03 45.22 61.79 86.18 64.11 0 30
AB3DMOT (IEEE IROS 2020)* [21] L 93.28 45.43 77.41 86.24 78.43 0 15
ACK3DMOT (IEEE T-IV 2023)* [39] L $N/A$ $N/A$ $N/A$ 88.73 86.81 8 68
ConvUKF (IEEE T-IV 2024)* [55] L 93.32 45.46 78.09 $N/A$ $N/A$ $N/A$ 17
FGO-based3DMOT (IEEE S-J 2024)* [56] L 93.77 46.14 77.85 86.53 79.00 1 $N/A$
GNN3DMOT (IEEE CVPR 2020)* [52] C+L 93.68 45.27 78.10 84.70 79.03 0 10
LEGO (Ours)* L 94.90 47.78 86.97 91.36 86.70 1 4
PolarMOT (ECCV 2022)** [42] L 94.32 $N/A$ $N/A$ 93.93 $N/A$ 31 $N/A$
CenterTube (IEEE T-MM 2023)*** [41] L 93.89 46.24 80.23 $N/A$ $N/A$ 78 $N/A$
EagerMOT (IEEE ICRA 2021)** [58] C+L 94.94 48.80 80.40 96.61 80.00 2 $N/A$
HIDMOT (IEEE T-VT 2023)** [64] C+L $N/A$ 45.64 79.68 90.45 81.44 $N/A$ $N/A$
CAMO-MOT (IEEE T-ITS 2023)** [9] C+L 95.20 48.04 81.48 $N/A$ $N/A$ $N/A$ $N/A$
Qiao et al., (IEEE IOT-J 2024)** [66] C+L $N/A$ 48.35 79.90 86.33 79.42 1 $N/A$
LEGO (Ours)** L 95.20 48.10 87.05 92.00 86.69 1 5
    The metrics are reported by using 3D clustering approach rather than any learning-based model as 3D object detector, according to [40].
    The metrics are reported by using PointRCNN [13] as 3D object detector.
    The metrics are reported by using PointGNN [14] as 3D object detector.
    The metrics are reported by using CenterPoint [15] as 3D object detector, which has much better detection performance than PointGNN in general.
TABLE IV: 3D MOT tracking results of the proposed method and other state-of-the-art trackers on Waymo [80] validation dataset.
[…]
SimTrack (ICCV 2021)+ [25] L 53.1 49.6 17.4 17.4
3DMODT (IEEE ICRA 2023)+ [81] L 55.9 51.2 18.9 18.9
SimpleTrack (ECCV Workshop 2022)* [28] L $N/A$ 56.9 $N/A$ $N/A$
Waymo LEGO (Ours)* L 58.30 58.30 19.44 19.44
PC3T (IEEE T-ITS 2022)** [45] L 61.3 61.3 $N/A$ $N/A$
Fast-Poly (IEEE RA-L 2024)** [82] L 62.3 62.3 $N/A$ $N/A$
LEGO (Ours)** L 63.1 63.1 23.6 23.6
    The metrics are reported without using pre-trained 3D object detector, but a joint detection and tracking network delicately for this task.
    The metrics are reported by using CenterPoint [15] as 3D object detector.
    The metrics are reported by using CasA [16] as 3D object detector.
IV-A Dataset and Evaluation Metrics
In this work, we use the KITTI dataset [78] and Waymo dataset [79] as many LiDAR-based MOT methods have been evaluated within these datasets. The efficacy of our proposed LEGO modular tracker is assessed through the Higher-Order Tracking Accuracy (HOTA) [77], defined as
$HOTA=\int_{0}^{1}HOTA_{\alpha}d\alpha\approx\frac{1}{19}\sum_{\alpha}HOTA_{\alpha},$ (20)
$HOTA_{\alpha}=\sqrt{\frac{\sum_{c}\mathcal{A}(c)}{FN+FP+TP}},$ (21)
[…]
$MOTA=1-\frac{\sum_{t}FN_{t}+FP_{t}+IDS_{t}}{\sum_{t}GT_{t}},$ (22)
where $FN_{t}$, $FP_{t}$, $IDS_{t}$, and $GT_{t}$ represent the number of false negatives, false positives, ID switch, and ground truth at time $t$ respectively. The MOTP is defined as
$MOTP=\frac{\sum_{t}dis_{t}}{\sum_{t}c_{t}},$ (23)
where $dis_{t}$ represents the distance between detection and its corresponding ground truth, and $c_{t}$ is the number of matched pairs at time $t$. Considering that the evaluation metrics on the KITTI benchmark are predominantly oriented towards a 2D perspective, additional comprehensive metrics pertinent to a 3D viewpoint are also employed, such as the Average Multiple Object Tracking Accuracy (AMOTA) and the corresponding precision metric, the Average Multiple Object Tracking Precision (AMOTP). AMOTA is defined as
$AMOTA=\frac{1}{M}\sum_{r}\left(1-\frac{\sum_{t}FN_{t}^{r}+FP_{t}^{r}+IDS_{t}^{r}}{\sum_{t}GT_{t}}\right),$ (24)
where $FN_{t}^{r}$, $FP_{t}^{r}$, $IDS_{t}^{r}$ represent the number of false negatives, false positives, ID switch at a specific recall value $r$ at time $t$, $M$ is the number of recall values respectively and AMOTP is defined as
$AMOTP=\frac{1}{M}\sum_{r}\left(1-MOTP_{r}\right),$ (25)
where $MOTP_{r}$ is the value of MOTP at a specific recall r{r}.
(a) Tracking ground truth for scene 0008
(b) PC3T tracking results for scene 0008
(c) LEGO tracking results for scene 0008
Figure 4: Visualization of 3D MOT results on the KITTI benchmark (scene 0008). The first row shows the ground truth tracking annotations across three consecutive frames. Each detected object in the current frame is represented by a bounding box along with its unique tracking ID. Objects are consistently colored according to their tracking IDS across frames. The second and third rows illustrate the tracking results obtained by PC3T and our proposed LEGO model, respectively, on the same frames. Please zoom in to see the details.
IV-B Implementation Details
During the training, Adam with a momentum set as 0.09 and an initial learning rate of 0.1 is applied. The learning rate decay is 0.001, and the training is executed across 100 epochs. During the inference phase, the LiDAR scanning frequency is fixed at 10Hz, and the threshold for the existence probability provided by the object detector is set to 0. This means that all the objects provided by the object detector are kept as input to the tracker. The track management threshold $N_{t}$ and $M_{t}$ are set to 14 and 4, respectively. Implementation details of the key modules are as follows:
Offset correction module: The kernel size, stride, and padding of the CNN layer are set to (3, 1, 1), respectively. The sizes of the two-layer MLP are set to (64, 6).
Feature extraction module: This module commences by setting the convolution channels to [64, 128, 1024] and configuring the kernel size and stride as 3 and 1, respectively, within the T-Net. Subsequently, a three-layer MLP is structured with sizes of (1024, 256, 128).
AMGN score calculation module: The number of AMGN blocks, $K$, is set to 3.
Time Complexity: Our method achieves a frame rate of 27.83 FPS on a single NVIDIA RTX 3080 GPU with an Intel(R) Xeon(R) Platinum CPU, measured across the entire KITTI validation dataset. Although direct FPS comparisons with existing methods are challenging due to differences in hardware configurations, our tracking pipeline requires no post-processing operations, thus functioning fully online and in real-time. Compared to previously reported methods such as FANTrack [19] (25.0 FPS), 3DT [72] (33.3 FPS), mmMOT [20] (4.8 FPS), and GNN3DMOT [52] (5.2 FPS). Although the hardware configurations used by these methods were not disclosed, limiting direct comparisons, our method achieves 27.83 FPS, which meets the requirements for online tracking speed.
IV-C Performance Comparison with Other State-of-the-Arts
In this subsection, the framework is evaluated and discussed concerning various MOT metrics.
IV-C1 Quality of Detection Input
The performance of a tracker is inherently tied to the effectiveness of the integrated detector. Three distinct object detectors were assessed in the conducted experiments: the CasA 3D detector, the PointGNN 3D detector, and the PointRCNN 3D detector. These 3D detectors were chosen based on their performance in the KITTI 3D Object detection challenge and arranged in descending order of effectiveness.
IV-C2 Performance Comparison with State-of-the-Art Trackers Using LiDAR Only on KITTI dataset with 2D MOT Metrics
The typical tracking result of our LEGO has been visualized in Fig. 4. The first row and the second show the ground truth and tracking result reported by our LEGO in scene 0006, the track state for every object in the current frame is marked by a bounding box and its track ID. The third row and the fourth-row show results in scene 0008. As delineated in Table I, a series of comparative analyses were conducted on various tracking methodologies within the context of the KITTI tracking benchmark. First, we compare our method with PC3T [45], utilizing PointRCNN as the object detector. The results reveal that our method achieved an enhancement of 0.25 in HOTA and 0.37 in MOTA on the testing dataset. Further analysis is extended to other methodologies utilizing disparate detectors, as outlined in Table I. For instance, in a comparison with CenterTube [41], which employs the same detector PointGNN, our method demonstrates an improvement of 8.32 in HOTA and 1.17 in MOTA. Likewise, when compared with UG3DMOT [43] that utilizes the CasA detector, our method achieved an enhancement of 2.15 in HOTA and 2.63 in MOTA.
IV-C3 Performance Comparison with State-of-the-Art Trackers Using LiDAR and Camera on KITTI dataset with 2D MOT Metrics
Many tracking algorithms leverage the fusion of 2D camera images and 3D point clouds to optimize performance within the KITTI tracking benchmark. Our proposed method, LEGO, was rigorously evaluated against a variety of LiDAR and camera fusion-based trackers. The comparative analysis was performed in two key segments in Table II, based on the detectors employed. Firstly, LEGO was compared with trackers utilizing PointRCNN as their detector. In this context, a notable improvement was observed against StrongFusionMOT [62], with LEGO registering 2.4 increments in HOTA and 3.44 increments in MOTA. The second segment of the comparative analysis involved trackers that employ PointGNN as their detector. LEGO’s improvements are again manifest, with a 3.62 enhancement in HOTA over HIDMOT [64] and a 5.34 advancement in HOTA and 0.09 advancement in MOTA over DualTracker [60]. A specific comparison with IMSF-MOT [57] highlighted LEGO’s superior performance.
IV-C4 Comparison of Proposed Method and Other State-of-the-Art Trackers on KITTI dataset with 3D MOT Metrics
Our proposed LEGO tracker has also been subjected to extensive comparative evaluation against various trackers for tracking in 3D space, by employing 3D MOT metrics, such as sAMOTA, AMOTA, and AMOTP. The details of these comparisons are tabulated in Table III. By employing PointRCNN as 3D detector, LEGO demonstrated the best overall performance across several fronts, comparing to all other recent proposed trackers using LiDAR only. Additionally, LEGO was evaluated against CenterTube [41] which utilizes CenterPoint as its 3D detector, generally considered to offer superior detection performance over PointGNN. Despite using PointGNN as the 3D detector, LEGO still managed to secure improvements of 1.31 in sAMOTA, 1.86 in AMOTA, and 6.82 in AMOTP. The comparisons were further extended to trackers with fusion of LiDAR and camera. Utilizing PointGNN as 3D detector, when juxtaposed with CAMO-MOT [9], LEGO matched the sAMOTA, but exhibited improvement in AMOTA and AMOTP, by 0.06 and 5.57, respectively. Besides, the LEGO system also demonstrates significant enhancements when compared with the recently proposed approach [66], achieving a 7.15 improvement in AMOTP and a 5.67 increase in MOTA.
IV-C5 Comparison of Proposed Method and Other State-of-the-Art Trackers on Waymo Dataset with 3D MOT Metrics
To demonstrate the generalization capability of our proposed LEGO tracker, the evaluation of LEGO tracker against other state-of-the-art methods on the Waymo dataset is also shown in Table IV. Using CenterPoint as the detector, LEGO demonstrates superior performance, achieving 58.30 MOTA for both L1 and L2 metrics. This represents a notable improvement of 1.4 over SimpleTrack on L2 MOTA. Furthermore, LEGO significantly outperforms previous methods, surpassing SimTrack and 3DMODT by 5.2 and 2.4 respectively on MOTA L1. LEGO also exhibits better performance with MOTP scores of 19.44 for both L1 and L2 metrics, demonstrating improvements of 2.04 and 0.54 compared to SimTrack and 3DMODT, respectively. When integrated with the CasA detector, LEGO’s performance improves with 63.1 MOTA and 23.6 MOTP.
IV-C6 Additional Analysis
Although our proposed LEGO tracker achieves superior HOTA and MOTA scores compared to other state-of-the-art methods (as shown in Table I), we observe an increased number of identity switches (IDS). The AMGN module primarily focuses on relational features derived from the current and immediately preceding frames. Without explicit long-term memory or robust temporal smoothing mechanisms, the tracker may face difficulties in maintaining consistent object identities over extended periods, especially within highly dynamic scenes, further increasing IDS. Moreover, the current absence of appearance-based features limits the stability of associations, exacerbating IDS issues. In future work, these limitations are planned to be addressed by incorporating a lightweight module or knowledge distillation integrated with appearance features, aiming to enhance the stability and robustness of associations and thereby reduce IDS.
IV-D Ablation Study for LEGO Modular Tracker
This section outlines the ablation study for various key modules in LEGO, all the experiments are implemented on KITTI car validation dataset.
IV-D1 Effectiveness of Modules
[…]
IV-D2 Effectiveness of Weight between Matrix $\bm{A}$ and Matrix $\bm{B}$
Throughout the experimental phase, refinements were introduced to the cost weight $w_{B}$ to ascertain the optimal equilibrium between the geometry and motion cost matrix $\bm{A}$, and the AMGN score matrix $\bm{B}$. A comprehensive assessment was conducted across four distinct levels for $w_{B}$, specifically 0, 1, 2, 3, and 4, as described in Table VI. When $w_{B}$ is set to 0, the AMGN module is disabled, meaning the graph structure and associated learned relationships are not utilized. Consequently, this configuration yields the lowest tracking performance. It was found that when $w_{2}$ was set to 2, the best outcomes were achieved. These results underscore the pivotal role played by the cost weight in determining the relative contributions of matrix $\bm{A}$ and matrix $\bm{B}$ to the association process.
IV-D3 Effectiveness of Parameter $K$ in AMGN Module
[…]
TABLE V: Ablation study of proposed modules on the KITTI car validation dataset.
[…]
82.600 85.01 88.00
83.079 85.954 90.649
84.800 87.20 91.00
85.808 88.63 92.65
TABLE VI: Ablation study of the weight $w_{B}$ for matrix $\bm{B}$ with proposed method on KITTI car validation dataset.
[…]
0 83.079 85.954 90.649 89.271
1 83.833 88.162 92.504 85.452
2 85.808 88.63 92.65 90.321
3 84.554 88.609 92.472 86.657
4 82.242 84.19 92.492 86.418
TABLE VII: Ablation study on the parameter K in AMGN Module with proposed method on KITTI car validation dataset.
[…]
0 82.600 85.01 88.00
1 83.357 86.592 90.677
2 84.263 87.014 91.588
3 85.808 88.63 92.65
4 85.189 88.038 92.257
V Conclusion
In this paper, we propose an online LiDAR-based tracker, LEGO, that introduces the offset correction module and AMGN score calculation module, and effectively integrates them into the existing tracking framework to tackle the inherent challenges of MOT. Our offset correction module demonstrates the capability to rectify certain errors in the detection results, contributing to improved tracking performance. Furthermore, the AMGN blocks which use the learnable adjacency matrix to estimate the association relationships between detected objects and predicted objects in an efficient way. However, it is important to acknowledge that the use of graph structures in our approach may lead to longer training times due to increased computational demands. As a potential solution for future research, techniques, such as voxel downsampling could be implemented to reduce the parameter count in the GNN, thereby enhancing computational efficiency and reducing time consumption.